OpenStudy (anonymous):

I am being asked a two part question and I cold use some help please....A) Write the expresion 27-2/3 in radical notation. B) Evaluate the radical expression. Please help

OpenStudy (anonymous):

Help!!!!!!!

OpenStudy (anonymous):

I am thinking that it could possibly be 3sq(2x-1)1/2\[3\sqrt{(2-1)1/2}\]

OpenStudy (anonymous):

Hello will someone please help me.... Please

OpenStudy (nottim):

wats radical notation?

OpenStudy (whpalmer4):

\[27^{-2/3} = \frac{1}{27^{2/3}} = \frac{1}{(\sqrt[3]{27)}^2} = \frac{1}{(3)^2} = \frac{1}{9}\]

OpenStudy (whpalmer4):

Here are the techniques I used: \[x^{-a} = \frac{1}{x^a}\] \[x^{a/b} = (\sqrt[b]{x})^a\] And of course 27 = 3 * 3 * 3 Does that all make sense?